A charged particle of mass 0.003g is held stationary in space by placing it in a downward direction of electric field of 6x10^4 N/C. Then, the magnitude of charge is?

Grade:12

1 Answers

Ankit Kumar
3 years ago
m=0.003g=3×10-6
E=6×104N/C
mg=F
(F=qE) 
mg=qE
3×10-6×10=q×6×104
By dividing the value of E by Left Hand Side you will get the answer written below
q=5×10-10C